OpenStudy (anonymous):

For the function f(x) = –(x + 1)2 + 4, identify the vertex, domain, and range.

OpenStudy (anonymous):

The vertex is (–1, 4), the domain is all real numbers, and the range is y ≥ 4. The vertex is (–1, 4), the domain is all real numbers, and the range is y ≤ 4.-my answer The vertex is (1, 4), the domain is all real numbers, and the range is y ≥ 4. The vertex is (1, 4), the domain is all real numbers, and the range is y ≤ 4.

OpenStudy (anonymous):

need help want to make sure if im correct

OpenStudy (anonymous):

ok what are you thinking?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

thank you so much can help make sure I have 3more problems correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the equation of the following graph in vertex form? parabolic function going down from the left through the point negative four comma zero and turning at the point negative three comma negative one and going up through the point negative two comma zero and then through the point zero comma eight and continuing up towards infinity

OpenStudy (anonymous):

Courtesy of Texas Instruments y = (x − 3)2 − 1 y = (x + 3)2 − 1 y = (x − 4)2 − 2-my answer y = (x − 4)2 + 8

OpenStudy (anonymous):

next time just type the numbers please. it's a lot easier to decipher. If it turns at (-3, -1) that's the vertex, so it should be the 2nd choice

OpenStudy (anonymous):

im so sorry

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Given the function f(x) = x2 and k = –1, which of the following represents a function opening downward?

OpenStudy (anonymous):

f(x) + k-my answer kf(x) f(x + k) f(kx)

OpenStudy (anonymous):

f(x)+k moves the function down k units. You want the one that makes -1*f(x)

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

b

OpenStudy (anonymous):

MMM o ok

OpenStudy (anonymous):

0 1 –1 -my answer –2

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

alculate the average rate of change for the given graph from x = –2 to x = 0

OpenStudy (anonymous):

rate of change just means slope. \[m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

plug the two points into the equation

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

well its between -2 nd 0

OpenStudy (anonymous):